# Multimodal machine learning with large language embedding model for polymer property prediction

### v1

@@ Abstract
Contemporary large language models (LLMs), such as GPT-4 and Llama, have harnessed extensive computational power and diverse text corpora to achieve remarkable proficiency in interpreting and generating domain-specific content, including materials science. To leverage the domain knowledge embedded within these models, we propose a simple yet effective multimodal architecture, PolyLLMem, which integrates text embeddings generated by Llama 3 with molecular structure embeddings derived from Uni-Mol, for polymer properties prediction tasks. In our model, Low-rank adaptation (LoRA) layers were also incorporated during the property prediction tasks to refine the embeddings based on our limited polymer dataset, thereby enhancing their chemical relevance for polymer SMILES representation. This balanced fusion of fine-tuned textual and structural information enables PolyLLMem to accurately predict a variety of polymer properties despite the scarcity of training data. Its performance is comparable to, and in some cases exceeds, that of graph-based models, as well as transformer-based models that typically require pretraining on millions of polymer samples. These findings demonstrate that LLM, such as Llama, can effectively capture chemical information encoded in polymer PSMILES, and underscore the efficacy of multimodal fusion of LLM embeddings and molecular structure embeddings in overcoming data scarcity and accelerating the discovery of advanced polymeric materials.
@@ 1 Introduction
The intrinsic complexity of polymer structures, coupled with the limited size of available databases, poses significant challenges for accurate property prediction. To overcome those limitations, researchers in polymer fields adapted a variety of strategies that encompass polymer representation, feature extraction, data augmentation, and application of diverse machine learning (ML) architectures to expedite the development of predictive models in polymer research. For instance, similar to Simplified Molecular-Input Line-Entry System (SMILES)9 used for small molecules, polymer SMILES including Polymer Simplified Molecular-Input Line-Entry System (PSMILES) and BigSMILES10, 11 were proposed as an extension of the traditional SMILES notation to describe macromolecular structures, including repeating units, end groups, and connectivity patterns. Furthermore, various molecular descriptors and structural representations, such as Morgan fingerprints, Morgan fingerprint with frequency, and molecular graphs12, 13, 14, 15, 16, 17, have been derived from polymers and utilized as input features for ML models. Those notations and feature extractions enable the standardized digital representation of polymers, facilitating computational analysis, database storage, and interoperability in polymer informatics applications. Additionally, to overcome the limitations imposed by the small size of polymer databases, Ma et al. developed the PI1M database by training a generative model on approximately 12,000 polymers manually collected from PolyInfo and subsequently generating around one million polymers18. Similarly, Kuenneth et al. generated a data set of 100 million hypothetical polymers by enumeratively combining chemical fragments extracted from a list of more than 13,000 synthesized polymers.11
With these advances in feature engineering, data augmentation, and database expansion, both classical ML and deep learning models have been employed to enhance polymer property prediction. Among classical ML models, ensemble tree-based models (Random Forests, Gradient Boosting) have emerged as top performers in polymer properties prediction tasks when provided with rich, chemically informed features.3, 19 Tao et al. conducted a comprehensive benchmark that evaluated 79 different ML models, including linear, tree-based, kernel, and deep learning approaches, on a large glass transition temperature ($T_{g}$) dataset using various polymer feature representations (e.g., Morgan fingerprints, RDKit descriptors, learned embeddings, and molecular graphs)20, 21. They demonstrated that a Random Forest model (RF) using a Morgan fingerprint with substructure frequency could achieve the highest accuracy in predicting $T_{g}$. Similarly, Uddin and Fan built an interpretable ML pipeline for $T_{g}$, testing nine algorithms (including decision trees, Support Vector Machine, k Nearest Neighbors, various boosting methods, and RF), and found that RF provided the highest predictive $R^{2}$.16 In addition to tree-based models, Wu et al. successfully applied Support Vector Regression and Gaussian Process to capture nonlinear structure–property relationships in polymers, emphasizing the importance of selecting robust descriptors for reliable predictions.14
On the deep learning front, architectures such as Graph Neural Networks (GNNs) have been employed to directly learn from polymer molecular graphs, capturing both local atomic environments and global connectivity patterns.22, 23, 24, 25 For example, Graph Convolutional Neural Networks (GCNNs) have been widely used to predict the dielectric constant and energy bandgap of polymers26, as well as $T_{g}$ and thermal decomposition temperature ($T_{d}$) of block copolymers and homopolymers27, 19. Gurnani et al. developed polyGNN, a model that learns directly from monomer structures to simultaneously predict multiple polymer properties.28 Similarly, transformer-based models, including ChemBERTa, TransPolymer, and polyBERT, have been pretrained on millions of PSMILES collections to capture rich contextual embeddings, which were subsequently applied to various downstream tasks.29, 11, 30, 31 TransPolymer integrated polymer-specific tokenization scheme to capture not only repeating units and end groups but also connectivity patterns, leading to high performance across a range of polymer properties prediction29. In addition, approaches such as polyBERT and PolyNC further enhance prediction performance by combining data augmentation, generative models, and multimodal fusion. PolyBERT utilized a massive dataset of 100 million hypothetical polymers to pretrain a chemical language model that can predict multiple properties simultaneously, while PolyNC fused natural language prompts with PSMILES to achieve robust multitask learning11, 15. Moreover, multimodal frameworks like MMPolymer 32 integrate 3D structural information with 1D sequence data, providing complementary insights that significantly improve predictions for properties sensitive to polymer conformation. Although these deep learning approaches demonstrated high accuracy in the predictions of polymer properties, they typically required careful architecture design and large volumes of real or virtual polymer data for pretraining before being effectively applied to downstream tasks.
@@
Recent advancements in machine learning, particularly with large language models (LLM) like GPT-4 and Llama, have leveraged vast computational resources and extensive text corpora, including materials science literature, to demonstrate remarkable capabilities in understanding, reasoning, and generating domain-specific content 33, 34, 35, 36, 37, 38. One study introduced GenePT, a model that leveraged LLM-derived embeddings from text descriptions of individual genes and single-cell data, and demonstrated performance comparable to or exceeding that of models pretrained on gene-expression profiles from millions of cells for downstream tasks such as gene-property and cell-type classification 39. Inspired by these results, we postulated that LLMs have already been exposed to substantial polymer-related knowledge and are capable of extracting relevant information from polymer texts for downstream tasks such as property prediction. Therefore, in this study, we leveraged an LLM to generate polymer-specific text embeddings. Moreover, to complement these textual embeddings and further enhance predictive capabilities, we incorporated structural embeddings obtained from Uni-Mol, a deep learning model specifically designed to encode detailed molecular structures 40. Uni-Mol has been pretrained on millions of 3D molecular representations of small molecules, enabling it to capture critical structural and chemical features necessary for chemical prediction tasks. By integrating embeddings from both the LLM and Uni-Mol, we developed PolyLLMem, a multimodal neural network tailored specifically for polymer property prediction. This approach substantially reduces the effort required to design and pretrain polymer-specific language models while also overcoming the challenge of limited experimental polymer datasets. In PolyLLMem, the complementary information from both textual and structural domains was balanced, and a Low-rank adaptation (LoRA) layer was incorporated to fine-tune the embeddings with our target small polymer dataset during the property prediction tasks. We evaluated PolyLLMem on predictions of 22 polymer properties, and our results revealed that the integrated approach yields performance comparable to, and in some cases exceeds, that of graph-based models, as well as transformer-based models that typically require pretraining on millions of polymer samples. The PolyLLMem offers several advantages: (i) it delivers robust performance on a wide range of property prediction tasks even when trained on a small dataset; (ii) it does not require extensive dataset curation, processing, or additional pretraining on the polymer data; and (iii) it is both fast to train and straightforward to implement.
@@ 2 Method
Figure 1: Schematic representation of PolyLLMem architecture. The architecture processes two inputs: text-based PSMILES string is encoded using the Llama 3 model to generate embeddings, while 3D molecular representations are processed by Uni-Mol to extract structural embeddings. These embeddings are merged after a LoRA layer to form a unified representation that is then employed in a single-task framework with a Multilayer Perceptron (MLP) for training and predicting polymer properties.
2.1 Data Collection
@@
2.2 Model Architecture
@@
Our multimodal model (PolyLLMem) integrated LLM-based and Uni-Mol-based embeddings to predict polymer properties by capturing both textual and structural information. The LLM-based embeddings were generated using the LLM Llama 3 by mean pooling the final hidden states (token-level embeddings) to produce a single embedding vector with 4096 dimensions for each input textual string34. Each input string, as shown in Figure. 1 follows the format “Polymer Smile: PSMILES,” where PSMILES represent the polymer SMILES. The mean pooling of the token-level embeddings effectively distills the rich chemical context captured by the LLM-based model into a robust representation.
@@
In parallel, we employed Uni-Mol embeddings, which were 1536-dimensional, to capture the 3D geometry and conformational details of the molecules. This approach yielded embeddings that encapsulate critical geometric relationships, providing complementary structural insights to the text-based representations obtained from the Llama 3. Moreover, since Uni-Mol does not recognize PSMILES, we replaced the asterisk * with "C" in the input for Uni-Mol. Once both embeddings were obtained, each was projected into a common latent space with a predefined hidden size using a linear layer with Gaussian Error Linear Unit (GELU) activation and batch normalization. LoRA layers further refined these projections, and a gated fusion mechanism dynamically combines the updated LLM and Uni-Mol representations.48 The fused embeddings are then processed through a refinement block and a regression network, which ultimately predicts the target polymer property. In our work, we focused on a single-property prediction task.
2.3 Training
@@
Our training procedure employed 5-fold cross-validation to ensure a robust evaluation of the model’s predictive performance. Training was performed using the AdamW optimizer with a learning rate scheduler (ReduceLROnPlateau) and an early stopping mechanism was applied to prevent overfitting.49, 50 Multiple loss functions, including Mean Square Error (MSE), Mean Absolute Error (MAE), and Huber, were used to guide optimization. For each fold, the best-performing model checkpoint was saved based on the validation loss, and final performance metrics (MAE and $R^{2}$) were computed on the test set and averaged across folds. Additionally, a grid search was used for hyperparameter tuning, optimizing key parameters such as hidden size, batch size, dropout rate, rank, alpha, learning rate, and weight decay.
@@
For baseline comparisons, we evaluated a suite of classical ML models, including RF, Linear Regression, Support Vector Regression(SVR), Decision Tree, Ridge Regression, AdaBoost, XGBoost, and a multilayer perceptron (MLP), using two distinct sets of input features separately: molecular descriptors (200 computed molecular properties) extracted from RDKit’s Descriptors.descList module, and Morgan fingerprints (MF) obtained from the RDKit package.20 Additionally, for further comparison, embeddings generated by the Llama 3 and Uni-Mol were separately evaluated using classical ML models.
@@ 3 Discussions
Figure 2: Two-dimensional Uniform Manifold Approximation and Projection (UMAP) plots of the embeddings generated from Llama 3 across all the polymers in this study. Panels (a–d) display colored dots representing property values for glass transition temperature ($T_{g}$), thermal decomposition temperature ($T_{d}$), chain band gap ($E_{gc}$), and Young’s modulus ($E$), respectively. Light gray dots indicate polymers with unknown property values.
@@
To evaluate the feasibility of using the LLM embedding model for polymer property predictions, we employed two-dimensional Uniform Manifold Approximation and Projection (UMAP) 51 to visualize the generated embeddings for all polymers in this study (see Figure 2). These embeddings were obtained by mean pooling the final layer of the Llama 3. In the UMAP plots, colored dots represent polymers with known property values for $T_{g}$ (Figure 2a), $T_{d}$ (Figure 2b), $E_{gc}$ (Figure 2c) and $E$ (Figure 2d), while light gray dots indicate polymers with unknown property values. In each plot, polymers with similar property values tend to form localized clusters of similar colors, with the exception of the $E$, where the differentiation was less distinct compared to the other properties. Nonetheless, this observation is noteworthy as it suggests that the LLM has successfully retained key chemical information and relationships inherent in the PSMILES strings, even though its predictive performance for properties like $E$ may not be as robust as for others.
Figure 3: Cosine similarity was computed for the token-level embeddings of the selected polymers (a–d): [*]CC([*])C, [*]CC([*])OCCCCCCCCCC, [*]CC([*])c1ccncc1, [*]CC([*])(F)C(=O)OCC(F)(F)C(F)(F)F. The resulting heatmaps display pairwise token similarities, with values near 1 indicating that tokens exhibit highly similar embedding representations, suggesting shared chemical or structural features.
@@
To further explore the LLM’s ability to capture chemical information in polymers, we extracted token-level embeddings. Rather than using the embeddings after mean pooling for UMAP visualization, we retained the individual token representations prior to mean pooling to compute the cosine similarity across all tokens for the selected polymer. In all the panel Figure 3a-d, tokens used for polymer notation such as parentheses (, ) or bracket placeholders [*], tend to have moderate similarity scores with each other, reflecting the model’s understanding of how they are used in PSMILES to denote branching, repeating units, or unspecified substituents. In contrast, lower similarity scores appear between tokens that represent clearly different chemical entities or notational functions (e.g., tokens related to side chains vs.[*] tokens). This separation indicates that the embeddings capture meaningful distinctions in chemical context.
@@
In Figure 3b, tokens corresponding to related chemical motifs, such as CCCC versus CCCCCC, show higher similarity scores, suggesting that the model recognizes these tokens as playing similar structural roles. Moreover, even among purely carbon tokens, subtle differences in length can lead to slightly different similarity scores. In Figure 3c, the PSMILES fragment consists of tokens such as c, 1, cc, n, cc, and 1, which collectively form a pyridine ring (e.g., c1ccncc1). These tokens show high inter-token similarity, indicating that the model recognizes them as parts of a unified aromatic ring structure and effectively captures their ring connectivity. In Figure 3d, the presence of clear clustering for fluorine tokens, along with the separation from purely carbon-based tokens, demonstrates the LLM’s ability to encode chemically relevant distinctions. Fluorine substituents have well-known effects on polymer properties (e.g., polarity, thermal stability), and the LLM embeddings’ separation of F from other tokens implies an internalized understanding of this difference. Overall, although the tokenizer used to train Llama may not perfectly differentiate certain aspects of polymer PSMILES (for instance, the tokens for [] and () are not distinctly separated), the LLM model still captures both syntactic and chemical relationships within PSMILES strings. While token-level embeddings retain considerably more information than mean-pooled embeddings, we opted for the aggregated embeddings for their simplicity and to demonstrate the effectiveness of the LLM embedding model in property prediction tasks.

### v2

@@ Abstract
Contemporary large language models (LLMs), such as GPT-4 and Llama, have harnessed extensive computational power and diverse text corpora to achieve remarkable proficiency in interpreting and generating domain-specific content, including materials science. To leverage the domain knowledge embedded within these models, we propose a simple yet effective multimodal architecture, PolyLLMem. By integrating text embeddings from Llama 3 with molecular structure embeddings from Uni-Mol, PolyLLMem enables the accurate prediction of polymer properties. Low-Rank Adaptation (LoRA) layers were integrated into our model during the property prediction stage to adapt the pretrained embeddings to our limited polymer dataset, thereby enhancing their chemical relevance for polymer SMILES representation. Such a balanced fusion of fine-tuned textual and structural information enables PolyLLMem to robustly predict a variety of polymer properties despite the scarcity of training data. Its performance is comparable to, and in some cases exceeds, that of graph-based or transformer-based models that typically require pretraining on millions of polymer samples. These findings demonstrate that LLM, such as Llama, can effectively capture chemical information encoded in polymer PSMILES, and underscore the efficacy of multimodal fusion of LLM embeddings and molecular structure embeddings in overcoming data scarcity and accelerating the discovery of advanced polymeric materials.
@@ 1 Introduction
Despite being intriguing, the intrinsic complexity of polymer structures, coupled with the limited size of available databases, poses significant challenges for accurate property prediction. To overcome these limitations, researchers in the polymer field have adopted a variety of strategies, including advancements in polymer representation, feature extraction, and data augmentation, to enhance the performance of diverse machine learning (ML) architectures and accelerate the development of predictive models in polymer research. For instance, similar to Simplified Molecular-Input Line-Entry System (SMILES)9 used for small molecules, polymer SMILES including Polymer Simplified Molecular-Input Line-Entry System (PSMILES) and BigSMILES10, 11 were proposed as an extension of the traditional SMILES notation to describe macromolecular structures, including repeating units, end groups, and connectivity patterns. Furthermore, various molecular descriptors and structural representations, such as Morgan fingerprint and its frequency-based variant, as well as molecular graphs 12, 13, 14, 15, 16, 17, have been derived from polymers and utilized as input features for ML models. Those notations and feature extractions enable the standardized digital representation of polymers, facilitating computational analysis, database storage, and interoperability in polymer informatics applications. Additionally, to overcome the limitations imposed by the small size of polymer databases, new datasets were generated. Examples included the PI1M database, constructed first by training a generative model on approximately 12,000 polymers manually collected from PolyInfo, subsequently generating around one million hypothetical polymers 18. Similarly, a dataset of 100 million hypothetical polymers was created by enumeratively combining chemical fragments extracted from over 13,000 synthesized ones.11
With these advances in data preparation, both classical ML and deep-learning models have been employed to enhance polymer property prediction. Classical models, such as ensemble tree-based methods, support vector regression, and Gaussian processes, have achieved fair performance in predicting properties on glass transition temperature ($T_{g}$) datasets when paired with chemically informed descriptors like Morgan fingerprints and RDKit features 3, 19, 20, 21, 16, 14.On the deep learning side, architectures including Graph Neural Networks (GNNs) and transformer-based models have been used to learn directly from polymer structures or PSMILES representations22, 23, 24, 25, 26, 27, 28. Recent models such as ChemBERTa, TransPolymer, and polyBERT leverage large-scale pretraining on polymer SMILES to generate contextual embeddings for downstream property prediction tasks29, 11, 30, 31. Other recent frameworks further enhance performance through multimodal fusion, integrating structural, textual, or generative information15, 32. Although these approaches demonstrated high accuracy in the predictions of polymer properties, they typically required careful architecture design and large volumes of real or virtual polymer data for pretraining before being effectively applied to downstream tasks.
@@
One of the most remarkable developments in machine learning is the emergence of large language models (LLMs), which have leveraged vast computational resources and extensive text corpora, including materials science literature, to demonstrate exceptional capabilities in understanding, reasoning, and generating domain-specific content.33, 34, 35, 36, 37, 38. Although originally developed for general-purpose natural language tasks, LLMs have shown strong potential in scientific domains when appropriately prompted or minimally tuned. For example, GenePT leveraged LLM-derived embeddings from gene descriptions and single-cell data, achieving performance comparable to or exceeding models pretrained on gene-expression profiles from millions of cells for tasks like gene-property and cell-type classification39. Similarly, LLMs, such as Llama, could also have been exposed to substantial polymer-related knowledge during pretraining and therefore, be capable of extracting relevant information from polymer texts for downstream tasks such as property prediction. Moreover, to complement the textual embeddings and further enhance predictive performance, we incorporated structural embeddings from Uni-Mol, a deep-learning model designed to encode detailed molecular structures40. Pretrained on millions of 3D molecular representations of small molecules, Uni-Mol could effectively capture essential chemical and spatial features relevant to prediction tasks.
By integrating embeddings from both the LLM and Uni-Mol, we developed PolyLLMem, a multimodal neural network tailored specifically for polymer property prediction. In our approach, the complementary information from both textual and structural domains was balanced, and a Low-rank adaptation (LoRA) layer was incorporated to fine-tune the embeddings with our target small polymer dataset during the property prediction tasks. We evaluated PolyLLMem on predictions of 22 polymer properties, and our results revealed that the integrated approach yields performance comparable to, and in some cases exceeds, that of graph-based or transformer-based models that typically require pretraining on millions of polymer samples. The PolyLLMem offers several advantages: (i) it demonstrates robust performance across a variety of property prediction tasks, even when trained on limited data; (ii) it requires minimal dataset curation, preprocessing, or additional pretraining on polymer-specific corpora; and (iii) it is computationally efficient and straightforward to implement, making it accessible for broad application.
@@ 2 Method
2.0.1 Data Collection
@@
2.0.2 Model Architecture
Figure 1: Schematic representation of PolyLLMem architecture. The architecture processes two inputs: text-based PSMILES string is encoded using the Llama3 model to generate embeddings, while 3D molecular representations are processed automatically by Uni-Mol to extract structural embeddings from a caped SMILES. These embeddings are merged after LoRA layers to form a unified representation that is subsequentially employed in a single-task framework with a Multilayer Perceptron (MLP) for training and predicting polymer properties.
@@ 2.0.2 Model Architecture
Our multimodal model (PolyLLMem) integrated LLM-based and Uni-Mol-based embeddings to predict polymer properties by capturing both textual and structural information. The LLM-based embeddings were generated using the LLM Llama3 by mean pooling the final hidden states (token-level embeddings) to produce a single embedding vector with 4096 dimensions for each input textual string34. Each input string, as shown in Figure. 1 follows the format “Polymer Smile: PSMILES.” (Figure1) The mean pooling of the token-level embeddings effectively distills the rich chemical context captured by the LLM-based model into a robust representation.
@@
In parallel, we employed Uni-Mol embeddings, which were 1536-dimensional, to capture the 3D geometry and conformational details of the molecules. This approach yielded embeddings that encapsulate critical geometric relationships, providing complementary structural insights to the text-based representations obtained from the Llama3. Moreover, since Uni-Mol does not recognize PSMILES, we replaced the asterisk * with "C" in the input for Uni-Mol (caped SMILES, Figure 1 ). Once both embeddings were obtained, each was projected into a common latent space with a predefined hidden size using a linear layer with Gaussian Error Linear Unit (GELU) activation and batch normalization. LoRA layers further refine these projections, and a gated fusion mechanism dynamically combines the updated LLM and Uni-Mol representations.48. The fused embeddings were subsequently processed through a refinement block and a dedicated regression network, where each network was responsible for predicting a single target property.
2.0.3 Training
@@
Our training procedure employed 5-fold cross-validation to ensure a robust evaluation of the model’s predictive performance. Training was performed using a weight-decay–regularized optimizer alongside a learning rate schedule that adaptively reduces the step size when validation performance plateaus. An early stopping mechanism was applied to prevent overfitting.49, 50 Multiple loss functions, including Mean Square Error (MSE), Mean Absolute Error (MAE), and Huber, were used to guide optimization. For each fold, the best-performing model checkpoint was saved based on the validation loss, and final performance metrics (MAE and $R^{2}$) were computed on the test set and averaged across folds. Additionally, a grid search was used for hyperparameter tuning, optimizing key parameters such as hidden size, batch size, dropout rate, rank, alpha, learning rate, and weight decay.
@@
For baseline comparisons, we evaluated a suite of classical ML models, including Random Forest (RF), Linear Regression (LR), Support Vector Regression (SVR), Decision Tree (DT), Ridge Regression (RR), AdaBoost, XGBoost, and a multilayer perceptron (MLP). Comparison was done using two distinct sets of input features separately: molecular descriptors (200 computed molecular properties) and Morgan fingerprints (MF) obtained from the RDKit package.20. Additionally, embeddings generated by the Llama3 and Uni-Mol were separately evaluated using classical ML models to understand their contributions.
@@ 3 Discussions
3.0.1 Rich Information in Embeddings
Figure 2: Upper: Two-dimensional Uniform Manifold Approximation and Projection (UMAP) plots of Llama3-generated polymer embeddings. The embeddings were first reduced to 100 dimensions using principal component analysis (PCA), followed by UMAP to project them into two dimensions. Panels (a–c) display colored dots representing property values for glass transition temperature ($T_{g}$), chain band gap ($E_{gc}$), and Young’s modulus ($E$), respectively. Colors represent the value of the property, while light gray dots indicate polymers with missing values. In both (a) & (b), A clear clustering of similar colors can be observed in each case, indicating that the LLM embeddings already capture meaningful chemical distinctions related to these properties prior to any task-specific training. Lower: Distributions of available data for each property. The proportion of known values relative to the entire dataset is also indicated.
@@ 3.0.1 Rich Information in Embeddings
To evaluate the feasibility of using the LLM embedding model for polymer property predictions, we employed two-dimensional Uniform Manifold Approximation and Projection (UMAP) 51 to visualize the generated embeddings for all polymers in this study (see Figure 2). These embeddings were obtained by mean pooling the final layer of the Llama3. In the UMAP plots, colored dots represent polymers with known property values for $T_{g}$ (Figure 2a), $E_{gc}$ (Figure 2b), and $E$ (Figure 2c), while light gray dots indicate polymers with unknown property values. In each plot, polymers with similar property values tend to form localized clusters of similar colors, with the exception of the Young’s modulus (Figure 2c), where the differentiation is less distinct compared to the other properties. Nonetheless, this observation is noteworthy as it suggests that the Llama3 has successfully retained key chemical information and relationships inherent in the PSMILES strings, even though its predictive performance for properties like Young’s modulus (Figure 2c) may not be as robust as for others, potentially due to the low availability of these data.
Figure 3: Cosine similarity was computed between token-level embeddings for three representative polymers, with values close to 1 indicating high embedding similarity and potential shared chemical or structural features. Selected polymers (a–c): [*]CC([*])C, [*]CC([*])c1ccncc1, [*]CC([*])(F)C(=O)OCC(F)(F)C(F)(F)F. Each example includes the molecular structure, a heatmap showing the full pairwise similarity matrix, and a chord diagram emphasizing strong inter-token relationships. In the chord diagrams, edges are drawn for token pairs exceeding a similarity threshold of 0.5 (0.7 used in c, for clarity). Different colors of nodes are used to show token locations. Line width represents similarity strength: blue edges connect tokens of the same character, while red edges indicate tokens of various names.
@@
UMAP projections revealed that embeddings from the Llama3 encoded basic domain knowledge related to certain properties. However, to ensure this knowledge is generalizable, the embeddings must also capture underlying chemical features—such as symmetry, similarity, and structural relationships—beyond property-specific patterns. Here, we extracted token-level embeddings to verify such ability. Rather than using the embeddings after mean pooling for UMAP visualization, we retained the individual token representations prior to mean pooling to compute the cosine similarity across all tokens for the selected polymer. The approach began by tokenizing the PSMILES using a Llama3 tokenizer, originally built for natural language, which occasionally divides chemical notations into segments that do not inherently represent meaningful chemical substructures. For example, a chemical group like “[*]” might be split into separate tokens (e.g., “[” and “*]”), leading to dispersed embeddings that hinder straightforward interpretability. To address this, after obtaining the initial embedding output for each token achieved from Llama3 tokenizer, we implemented a custom embedding merging strategy where token representations that should collectively denote a single structural unit were averaged together. This involved selectively combining specific token embeddings, such as merging adjacent tokens or even portions of tokens, to realign the representation with the underlying polymer structure. Then the refined token level embeddings were used to calculate the cosine similarity among the refined tokens as shown in Figure 3. In all the panels in Figure 3a-c, tokens used for polymer notation[*], tend to have moderate similarity scores (0.7) with each other, reflecting the model’s understanding of how they are used in PSMILES to denote branching, repeating units, or unspecified substituents. In contrast, lower similarity scores appear between tokens that represent clearly different chemical entities or notational functions (e.g., tokens related to side chains vs.[*] tokens). This separation indicates that the embeddings capture meaningful distinctions in chemical context.
@@
In Figure 3a, the backbone token CC and the sidechain token C have significantly lower similarities, despite both being carbons. In Figure 3b, the PSMILES fragment consists of tokens such as c, 1, cc, n, cc, and 1, which collectively form a pyridine ring (e.g., c1ccncc1). These tokens show moderate inter-token similarity, indicating that the model recognizes them as parts of a unified aromatic ring structure and effectively captures their ring connectivity. The symmetry of molecular structure is also preserved, exemplified by the tokens of cc. In Figure 3c, the presence of clear clustering for fluorine tokens, along with the separation from purely carbon-based tokens, demonstrates the Llama3’s ability to encode chemically relevant distinctions. Fluorine substituents have well-known effects on polymer properties (e.g., polarity, thermal stability), and the LLM embeddings’ separation of F from other tokens implies an internalized understanding of this difference. Note that now, the tokens CC and C, both inside the sidechain, report higher similarity (0.8) than that of Figure 3a. Overall, although the tokenizer in Llama3 may not perfectly differentiate certain aspects of polymer PSMILES, our approach to refining the tokens shows that the Llama3 can still effectively capture both the syntactic structure and the underlying chemical relationships within the PSMILES strings. While token-level embeddings retain considerably more information than mean-pooled embeddings, we opted for the aggregated embeddings for their simplicity and to demonstrate the effectiveness of the LLM embedding model in property prediction tasks.
3.0.2 Performance of PolyLLMem